OpenStudy (anonymous):

write the equation in standard form. Is parallel to y= 3x-5 and passes through points (2, -3) ??

HanAkoSolo (jamierox4ev3r):

y=3x-5 -3=3(2)+b -3=6+b -6 -6 ----- b=-9

HanAkoSolo (jamierox4ev3r):

because the lines are parallel the slope are the same and your answer would be: y=3x-9